OpenStudy (anonymous):

what is the soultion sqrt5x+1 - sqrt x=5

OpenStudy (anonymous):

@hartnn @jjuden

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (studygurl14):

Is this the equation you need help with, @Jelia97 ? \[\sqrt{5x+1} -\sqrt{x} = 5\]

OpenStudy (anonymous):

yes thats it

OpenStudy (studygurl14):

Okay, so the first thing you want to do is get rid of the square roots, so you want to square both sides so you get \[(\sqrt{5x+1} - \sqrt{x})^{2} = 5^{2}\] \[5x + 1 - \sqrt{5xx+1} - \sqrt{5xx+1} + x\] What do you do now?

OpenStudy (studygurl14):

oops, I forgot to put = 25 at the end

hartnn (hartnn):

it would be better if you first add sqrtx on both sides, before squaring both sides

OpenStudy (studygurl14):

What's that @hartnn ?

OpenStudy (studygurl14):

sqrtx?

hartnn (hartnn):

yes like this sqrt(5x+1) - sqrt x=5 sqrt(5x+1) = sqrt x+5 now square both sides

OpenStudy (studygurl14):

Oh, I know what you mean now. Oh, yeah. I probably should have done that first.

OpenStudy (studygurl14):

\[\sqrt{5x + 1} - \sqrt{x} = 5\] \[\sqrt{5x + 1} = 5 + \sqrt{x}\] \[(\sqrt{5x + 1})^{2} = (\sqrt{x} + 5)\] \[5x + 1 = x + 5\sqrt{x} + 5\sqrt{x} + 25\] Okay, so NOW what do you do?

OpenStudy (anonymous):

is the answer x=16 and x=1

hartnn (hartnn):

16, yes. 1 no

OpenStudy (anonymous):

u sure?? is it just 16

hartnn (hartnn):

yup, only 16

OpenStudy (anonymous):

oh my mistake plugged in a extra 1 lol nvm

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^